### v3

# Borsuk-Ulam and Replicable Learning of Large-Margin Halfspaces

Abstract
Recent remarkable advances in learning theory have established that, for total concept classes, list replicability, global stability, differentially private (DP) learnability, and shared-randomness replicability all coincide with the finiteness of Littlestone dimension. Does this equivalence extend to partial concept classes?
We answer this question by proving that the list replicability number of $d$-dimensional $\gamma$-margin half-spaces satisfies
$\frac{d}{2}+1\leq\operatorname{LR}(\mathcal{H}^{d}_{\gamma})\leq d,$
which grows with dimension. Consequently, for partial classes, list replicability and global stability do not necessarily follow from bounded Littlestone dimension, pure DP-learnability, or shared-randomness replicability.
Applying our main theorem, we resolve several open problems:
  • Every disambiguation of infinite-dimensional large-margin half-spaces to a total concept class has unbounded Littlestone dimension, answering an open question of Alon, Hanneke, Holzman, and Moran (FOCS ’21).
  • The maximum list-replicability number of any finite set of points and homogeneous half-spaces in $d$-dimensional Euclidean space is $d$, resolving a problem of Chase, Moran, and Yehudayoff (FOCS ’23).
  • Every disambiguation of the Gap Hamming Distance problem in the large gap regime has unbounded public-coin randomized communication complexity. This answers an open question of Fang, Göös, Harms, and Hatami (STOC ’25).
  • There exists a partial concept class with Littlestone dimension $1$ such that all its disambiguations have infinite Littlestone dimension. This answers a question of Cheung, H. Hatami, P. Hatami, and Hosseini (ICALP ’23).
Our lower bound follows from a topological argument based on the local Borsuk-Ulam theorem of Chase, Chornomaz, Moran, and Yehudayoff (STOC ’24). For the upper bound, we construct a list-replicable learning rule using the generalization properties of SVMs.
1 Introduction
The large-margin half-space classification problem is a central topic in learning theory, extensively studied for its theoretical significance and practical impact. Theoretically, it is a fundamental and elegant geometric problem, and it serves as a simple model for understanding more complex learning problems. Practically, its importance stems from the success of Support Vector Machines (SVMs), which leverage the large-margin assumption to produce accurate classifications in high-dimensional spaces. SVMs play a key role in many real-world applications, including pattern and text recognition, image analysis, geographic information systems, bioinformatics, healthcare, and fraud detection.
To describe the large-margin classification problem, we first consider standard half-space classification without any margin assumption. After appropriate normalization and homogenization, we may assume that the domain is the unit sphere $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$, and the objective is to learn an unknown homogeneous half-space which classifies each point on the sphere as either ±$\pm 1$ depending on whether it belongs to the half-space.
The standard mathematical framework for analyzing the complexity of a learning task is probably approximately correct (PAC) learning. In this setting, the learner receives a training set of labeled points drawn independently from an unknown distribution $\mathcal{D}$, and the goal is to produce a hypothesis that accurately predicts the labels of new points sampled from $\mathcal{D}$ with high probability. The fundamental theorem of PAC learning states that the size of the training set required for PAC learning depends on a combinatorial parameter known as the VC dimension (Definition 2.5).
We also consider the Littlestone dimension (Definition 2.7), a refinement of the VC dimension that determines the optimal number of mistakes in online learning. In online learning, the learner receives data points sequentially and must predict the label of each point in turn before observing its correct label. The goal is to minimize the number of mistakes throughout the learning process.
Since the VC dimension of homogeneous half-spaces in $\mathbb{R}^{d}$ is $d$, PAC-learning this class requires a large training set in the high-dimensional settings. Moreover, the Littlestone dimension of this class is infinite except in the trivial case of $d=1$, meaning the class is not online learnable, even in $\mathbb{R}^{2}$.
Under the large-margin assumption, the learning task is restricted to points that lie at least a margin $\gamma>0$ away from the defining hyperplane. This constraint allows algorithms such as Perceptron [MP43, Ros58] and Support Vector Machines [VC64, CV95] to achieve efficient PAC and online learning, even in high-dimensional spaces. We study the large-margin halfspaces problem through the formal lens of partial classes, which offers a general framework for analyzing such constrained learning tasks.
A partial concept class over an arbitrary domain $\mathcal{X}$ is a set {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$, where each $h\in\mathcal{H}$ is called a partial concept. The value $h(x)=\star$ indicates that $h$ is undefined at $x$, and therefore, both ±$\pm 1$ are acceptable predictions for the label of $x$.
In the large-margin setting, the domain is $\mathbb{S}^{d-1}$ and every homogeneous half-space defines a partial concept that assigns $h(x)=\star$ if $x$ lies within distance $\gamma$ of the defining hyperplane of $h$. Otherwise, it classifies $x$ as ±$\pm 1$ depending on whether it belongs to the half-space. More formally, each hypothesis $h_{w}:\mathbb{S}^{d-1}\to\{\pm 1,\star\}$ is specified by a unit vector $w\in\mathbb{S}^{d-1}$ and given by
$h_{w}(x)\coloneqq\begin{cases}\operatorname{sgn}(\langle w,x\rangle)&\text{if }|\langle w,x\rangle|\geq\gamma\\ \star&\text{otherwise}\end{cases}.$ (1)
We denote the partial concept class of all such $h_{w}$ by $\mathcal{H}^{d}_{\gamma}$.
We also define $\mathcal{H}^{\infty}_{\gamma}$ as the class of all partial concepts $h_{w}:\bigcup_{d\in\mathbb{N}}\mathbb{S}^{d-1}\to\left\{\pm 1,\star\right\}$, where each hypothesis is specified by a unit vector $w\in\bigcup_{d\in\mathbb{N}}\mathbb{S}^{d-1}$ of arbitrary finite dimension. For any $x$ and $w$ of different dimensions, we set $h_{w}(x)\coloneqq\star$, and otherwise, define $h_{w}(x)$ as in Eq. 1.
In [AHHM22], Alon, Hanneke, Holzman, and Moran proved that, as in the case of total concept classes, the VC and Littlestone dimensions characterize PAC and online learnability of partial concept classes. In the case of $\mathcal{H}_{\gamma}^{d}$, the classic mistake-bound analysis of the Perceptron algorithm [MP43, Ros58] (see also [SSBD14, Theorem 9.1]) shows the following upper bound on the Littlestone and VC dimensions:
$\operatorname{VCdim}(\mathcal{H}_{\gamma}^{d})\leq\operatorname{Ldim}(\mathcal{H}_{\gamma}^{d})\leq\gamma^{-2}.$ (2)
Crucially, these bounds are independent of $d$, which explains the efficient PAC and online learnability of $\mathcal{H}_{\gamma}^{d}$ in arbitrarily high-dimensional spaces.
Recent advances in learning theory, sparked by the influential works on the differential privacy of PAC learning [ALMM19, BLM20, ABL+22], have established that for total concept classes, finite Littlestone dimension not only characterizes online learnability but is also equivalent to list replicability, global stability, differential privacy (DP) learnability, and shared-randomness replicability. We will discuss these notions in depth shortly. This naturally leads to the question of whether these equivalences extend to partial concept classes. The primary goal of this paper is to provide a negative answer to this question by proving that $\mathcal{H}_{\gamma}^{\infty}$ is not list replicable, even though it has a bounded Littlestone dimension by Eq. 2.
Before delving deeper into these results, we formally define some key notions. Consider a partial concept class {±$\mathcal{H}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$. A distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$ is realizable by a partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ if, for every $n$, a random sample $\bm{S}=((\bm{x}_{i},\bm{y}_{i}))_{i=1}^{n}\sim\mathcal{D}^{n}$ is almost surely realizable by some $h\in\mathcal{H}$, meaning that $h(\bm{x}_{i})=\bm{y}_{i}$ for all $i=1,\ldots,n$. Note that $\mathcal{D}$ is a distribution over $\mathcal{X}\times\left\{\pm 1\right\}$, so none of the labels $\bm{y}_{i}$ take the value $\star$.
Here, and throughout the paper, we use boldface letters to denote random variables and use the notation $(\bm{x},\bm{y})\sim\mathcal{D}$ to express that $(\bm{x},\bm{y})$ is a random variable distributed according to $\mathcal{D}$.
The population loss of a partial concept $h\in\left\{\pm 1,\star\right\}^{\mathcal{X}}$ with respect to a distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$ is
$\mathcal{L}_{\mathcal{D}}(h)\coloneqq\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}[h(\bm{x})\neq\bm{y}].$
Throughout this work, a learning rule refers to a (randomized) function $\bm{\mathcal{A}}$ that maps any sample $S\in\bigcup_{n=0}^{\infty}(\mathcal{X}\times\left\{\pm 1\right\})^{n}$ to a hypothesis $\bm{\mathcal{A}}(S)\in\left\{\pm 1\right\}^{\mathcal{X}}$. Since our primary focus is sample complexity rather than computational efficiency, we impose no computability constraints on $\bm{\mathcal{A}}$.
Replicability.
Replicability refers to the ability of an algorithm or study to produce consistent results when repeated under similar conditions and with similar data. It is a key principle of the scientific method that ensures research findings are reliable and not due to chance.
In recent years, a rapidly growing body of research has emerged that introduced various rigorous formulations of replicability for learning algorithms, particularly in the context of the PAC learning framework [BLM20, MM22, CMY23, BGH+23, KVYZ23, EKK+23, EKM+23, MSS23, EHKS23, KKL+24, KKMV23]. We introduce only the notions that are central to our main results and refer the reader to prior works, such as [MSS23, KKMV23, MM22], for a discussion of other related notions. Throughout, we use replicability as an umbrella term referring to the various formulations of this general concept.
The concept of replicability in PAC learning first emerged in [BLM20, ABL+22] in the study of differential privacy of PAC learning algorithms. These works introduced a notion of replicability known as global stability to derive privacy guarantees from online learnability.
A learning rule $\bm{\mathcal{A}}$ for a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is $(\epsilon,\rho)$-globally stable if for every realizable distribution $\mathcal{D}$, there is a hypothesis $h\in\left\{\pm 1\right\}^{\mathcal{X}}$ with population loss $\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon$ satisfying
$\Pr_{\bm{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\bm{S})=h]\geq\rho,\ \ \text{where }n=n(\epsilon).$
In other words, when run with samples from $\mathcal{D}$, there is a non-negligible chance $\rho>0$ that the learner will output the same hypothesis $h$. We define $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})$ to be the supremum of $\rho$ such that there is a $(\epsilon,\rho)$-globally stable learner for $\mathcal{H}$. The global stability parameter of $\mathcal{H}$ is then defined as
$\rho^{\mathrm{gs}}(\mathcal{H})\coloneqq\inf_{\epsilon>0}\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H}).$
The definition of global stability might initially seem weak, as a globally stable learner is not necessarily a PAC learner. In particular, since $\rho$ can be a small constant, there may be a probability as great as $1-\rho$ that the learning rule outputs a hypothesis with large population loss. However, [CMY23] showed that for total classes {±$\mathcal{H}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$, global stability is equivalent to a stronger notion called list replicability.
Definition 1.1 (List replicability).
A learning rule $\bm{\mathcal{A}}$ with sample complexity $n(\epsilon,\delta)$ is said to be an $(\epsilon,L)$-list replicable learner for {±$\mathcal{H}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$ if the following holds. For every realizable distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$, there is a list of total hypotheses $h_{1},\ldots,h_{L}\in\left\{\pm 1\right\}^{\mathcal{X}}$ such that
  • The population loss of every $h_{i}$ with respect to $\mathcal{D}$ is at most $\epsilon$:
    $\mathcal{L}_{\mathcal{D}}(h_{i})\leq\epsilon\text{ for every }i=1,\ldots,L.$
  • The probability that $\bm{\mathcal{A}}$ outputs a hypothesis outside the list is at most $\delta$:
    $\Pr_{\bm{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\bm{S})\not\in\left\{h_{1},\ldots,h_{L}\right\}]\leq\delta\ \text{ where }n=n(\epsilon,\delta).$
Given an accuracy parameter $\epsilon\in(0,1)$, the $\epsilon$-list replicability number of $\mathcal{H}$, denoted by $\operatorname{LR}_{\epsilon}(\mathcal{H})$, is the smallest integer $L$, such that there exists an $(\epsilon,L)$-list replicable learner for $\mathcal{H}$. We define $\operatorname{LR}_{\epsilon}(\mathcal{H})=\infty$ if no such integer exists.
The list replicability number of $\mathcal{H}$ is $\operatorname{LR}(\mathcal{H})\coloneqq\sup_{\epsilon>0}\operatorname{LR}_{\epsilon}(\mathcal{H}).$ We say $\mathcal{H}$ is list replicable if $\operatorname{LR}(\mathcal{H})<\infty$.
Definition 1.1 provides a strong notion of replicability as the learner’s output is typically chosen from a small list $\left\{h_{1},\ldots,h_{L}\right\}$, and all these hypotheses have small population loss.
In [CMY23], Chase, Moran and Yehudayoff proved that for every total class {±$\mathcal{H}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$, list replicability is equivalent to global stability. It is easy to check that their proof applies to partial concept classes, resulting in the following relationship between the list replicability number and global stability parameter.
Theorem 1.2 ([CMY23]).
For every partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$, we have
$\rho^{\mathrm{gs}}(\mathcal{H})=\frac{1}{\operatorname{LR}(\mathcal{H})}.$
For completeness, we provide a proof in Appendix A.
As shown in [BLM20], the global stability parameter of any nontrivial concept class is at most $1/2$. This fundamental limitation implies that no nontrivial class can achieve true replicability in the strongest sense, where the learning algorithm outputs the same hypothesis with high probability. To remedy this, Impagliazzo, Lei, Pitassi, and Sorrell [ILPS22] introduced a weaker notion of replicability where different executions of the algorithm can use the same random seed. Formally, let $\mathcal{A}(S,r)$ be a randomized learning rule, where $r$ denotes the random seed.
Definition 1.3 (Shared-randomness replicability).
A concept class {±,}$\mathcal{H}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$ is shared-randomness replicable if there exists a learning rule $\mathcal{A}$ and a sample complexity function $n(\epsilon,\delta)$ such that, for every $\epsilon,\delta>0$ and every realizable distribution $\mathcal{D}$, the following conditions hold:
  • Small population loss: $\Pr_{\bm{S}\sim\mathcal{D}^{n},\bm{r}}[\mathcal{L}_{\mathcal{D}}\left(\mathcal{A}(\bm{S},\bm{r})\right)>\epsilon]\leq\delta.$
  • Replicability with shared randomness: $\Pr_{\bm{S,S^{\prime}}\sim\mathcal{D}^{n},\bm{r}}[\mathcal{A}(\bm{S},\bm{r})=\mathcal{A}(\bm{S^{\prime}},\bm{r})]\geq 1-\delta.$
The following theorem, which brings together several celebrated recent results in learning theory, establishes that for total concept classes, all these notions of replicability, along with approximate differential privacy, are characterized by the finiteness of the Littlestone dimension.
Theorem 1.4 ([ALMM19, BLM20, ABL+22, CMY23, ILPS22]).
Let {±$\mathcal{H}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$ be a total concept class. The following statements are equivalent.
  • $\operatorname{Ldim}(\mathcal{H})<\infty$.
  • $\mathcal{H}$ is globally stable and equivalently list replicable.
  • $\mathcal{H}$ is shared-randomness replicable.
  • $\mathcal{H}$ is approximately differentially private (DP)-learnable (See Definition 2.9).
1.1 Main contributions
Replicability of large-margin half-spaces.
The differentially private and replicable learning of large-margin half-spaces have been studied extensively [BDMN05, LNUZ20, BMNS19, KMST20, BCS20, BMS22a, BMS22b, ILPS22, KKL+24].
In [LNUZ20], Nguyen, Ullman, and Zakynthinou gave a DP-learner for $\mathcal{H}^{d}_{\gamma}$ with a sample complexity that does not depend on the dimension $d$. In fact, they proved a stronger result that $\mathcal{H}^{d}_{\gamma}$ is pure DP-learnable (Definition 2.10). Combining the dimension-free DP-learnability of $\mathcal{H}^{d}_{\gamma}$ with the “DP-learnable to Shared-randomness Replicability” reduction of [BGH+23] shows that $\mathcal{H}^{d}_{\gamma}$ is shared-randomness replicable with dimension-free sample complexity. More recently, Kalavasis, Karbasi, Larsen, Velegkas, and Zhou [KKL+24] employed a more sophisticated rounding scheme to directly obtain a shared-randomness replicability of this problem, which further improves the sample complexity. In the following theorem, we summarize the consequences of these lines of work that are relevant to this paper.
Theorem 1.5 ([MP43, Ros58, LNUZ20, KKL+24]).
For every $\gamma>0$ and $d\in\mathbb{N}\cup\left\{\infty\right\}$, the class $\mathcal{H}^{d}_{\gamma}$ satisfies the following.
  • $\operatorname{Ldim}(\mathcal{H}^{d}_{\gamma})<\gamma^{-2}$.
  • The class $\mathcal{H}^{d}_{\gamma}$ is (pure) DP-learnable with dimension-independent sample complexity.
  • The class $\mathcal{H}^{d}_{\gamma}$ is shared-randomness replicable with dimension-independent sample complexity.
Given the positive results in Theorem 1.5 and the equivalences in Theorem 1.4, one might expect that, as in the case of total concepts, the list replicability number of $\mathcal{H}^{d}_{\gamma}$ is independent of $d$. Our main theorem establishes that this is not the case—the list replicability number of $\mathcal{H}^{d}_{\gamma}$ grows as $d$ increases.
Theorem 1.6 (Main Theorem).
For any fixed margin $\gamma\in(0,1)$, finite dimension $d>1$, and accuracy parameter $\epsilon\in(0,1/2)$,
$\frac{d}{2}+1\leq\operatorname{LR}_{\epsilon}(\mathcal{H}_{\gamma}^{d})\leq d.$
In particular, $\mathcal{H}_{\gamma}^{\infty}$ is not list replicable (equiv. globally stable).
Theorem 1.6 reveals a surprising distinction in the partial setting: list replicability does not follow from bounded Littlestone dimension, replicability, DP-learnability, or even pure DP-learnability. Specifically, for every $\gamma\in(0,1)$, the partial concept class $\mathcal{H}^{\infty}_{\gamma}$ is shared-randomness replicable, pure DP-learnable, and has a finite Littlestone dimension, but it is not list replicable.
The lower bound in Theorem 1.6 relies on a topological argument involving covers of the sphere by antipodal-free open sets. In particular, we apply the local Borsuk-Ulam theorem of [CCMY24], which states that in such a cover, there is a point that belongs to at least $\frac{d}{2}+1$ sets. Alternatively, one could use Ky Fan’s classical theorem [Fan52], but this would yield the slightly weaker bound of $\frac{d}{2}$.
For the upper bound, we construct a learning rule that uses the generalization properties of hard-SVM combined with a list-replicable rounding scheme using a fine net in general position.
Disambiguations of Large-margin Half-spaces:
One of the central questions studied in [AHHM22] is whether the learnability of a partial concept class can always be inferred from the learnability of a suitable extension of it to a total concept class.
Formally, a disambiguation of a partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is a total concept class ¯{±$\overline{\mathcal{H}}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$ such that for every $h\in\mathcal{H}$ and every finite $S\subseteq h^{-1}(\left\{\pm 1\right\})$, there exists an $\bar{h}\in\overline{\mathcal{H}}$ that is consistent with $h$ on $S$. Intuitively, one may think of disambiguation as replacing each $\star$ with $-1$ or $+1$, although this intuition is not completely rigorous in the infinite case.
Note that any distribution realizable by a partial concept class $\mathcal{H}$ is also realizable by any disambiguation $\overline{\mathcal{H}}$ of $\mathcal{H}$. Consequently, if a disambiguation $\overline{\mathcal{H}}$ is efficiently learnable, the same holds true for $\mathcal{H}$.
Motivated by the fact that $\operatorname{Ldim}(\mathcal{H}_{\gamma}^{d})\leq\gamma^{-2}$, [AHHM22] posed the following question.
Question 1.7 ([AHHM22, Question 4]).
Does there exist a disambiguation of $\mathcal{H}_{\gamma}^{d}$ by a total concept class whose VC/Littlestone dimension is bounded by a function of $\gamma$?
In the following theorem, we provide a negative answer to the Littlestone dimension part of 1.7.
Theorem 1.8.
For every $d\in\mathbb{N}$, every disambiguation $\overline{\mathcal{H}}$ of $\mathcal{H}_{\gamma}^{d}$ satisfies $\operatorname{Ldim}(\overline{\mathcal{H}})=\Omega(\sqrt{\log d})$.
Proof.
This follows by combining the lower bound of Theorem 1.6 and a result of [GGKM21] showing that for every total concept class $\mathcal{H}$, $\operatorname{LR}_{\epsilon}(\mathcal{H})\leq 2^{O_{\epsilon}\left(\operatorname{Ldim}(\mathcal{H})^{2}\right)}$. ∎
On the other hand, there is a disambiguation $\overline{\mathcal{H}}$ of $\mathcal{H}_{\gamma}^{d}$ with Littlestone dimension $\operatorname{Ldim}(\overline{\mathcal{H}})=O(d)$ since [HHM23] shows that there is a disambiguation of $\mathcal{H}_{\gamma}^{d}$ into a finite total concept class $\overline{\mathcal{H}}$ with $2^{O(d)}$ concepts. The upper bound on the Littlestone dimension follows from the fact that every finite class $\mathcal{H}$ satisfies $\operatorname{VCdim}(\mathcal{H})\leq\operatorname{Ldim}(\mathcal{H})\leq\log|\mathcal{H}|$.
Theorem 1.8 has two important implications. Firstly, it shows $\mathcal{H}^{\infty}_{\gamma}$ is a partial class with finite Littlestone dimension, but any disambiguation of it has an infinite Littlestone dimension. The existence of such partial classes was posed as a question in [AHHM22] and answered in [CHHH23] through a complex construction based on Göös’ breakthrough [Göö15] in communication complexity. Theorem 1.8 offers a more natural example of a concept class with this property.
Secondly, the construction of  [CHHH23] has Littlestone dimension $2$, and in [CHHH23, Question 4.1], they asked whether there exists a partial concept class with Littlestone dimension $1$ such that all its disambiguations have infinite Littlestone dimension. The following corollary of Theorem 1.8 resolves their question.
Corollary 1.9.
Let $\gamma\in(1/\sqrt{2},1)$. Then $\operatorname{Ldim}(\mathcal{H}_{\gamma}^{\infty})=1$, while any disambiguation of $\mathcal{H}_{\gamma}^{\infty}$ has infinite Littlestone dimension.
Proof.
We have $21\leq\operatorname{Ldim}(\mathcal{H}_{\gamma}^{\infty})\leq 1/\gamma^{2}<$ 2. An application of Theorem 1.8 completes the proof. ∎
List replicability of finite classes.
In [CMY23, Theorems 5 and 13], Chase, Moran, and Yehudayoff proved that the list replicability number of any finite set of homogeneous half-spaces and points that do not lie on the half-spaces in $\mathbb{R}^{2}$ is at most $2$. They further asked whether a similar finite bound holds in higher dimensions. As we discuss below, we give an affirmative answer to their question in Theorem 1.11.
Definition 1.10.
Define the finitary list replicability number of a concept class {±$\mathcal{H}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$, denoted $\widetilde{\operatorname{LR}}(\mathcal{H})$, as
$\widetilde{\operatorname{LR}}(\mathcal{H})\coloneqq\sup_{\text{finite }S\subseteq\mathcal{X}}\operatorname{LR}(\mathcal{H}|_{S}).$
Given $d\in\mathbb{N}$, let $\mathcal{H}^{d}$ be the partial concept class defined by points in $\mathbb{S}^{d-1}$ and partial concepts $h_{w}:\mathbb{S}^{d-1}\to\left\{\pm 1,\star\right\}$, where $w\in\mathbb{S}^{d-1}$, given by
$h_{w}(x)\coloneqq\begin{cases}\operatorname{sgn}(\langle w,x\rangle)&\text{if }\langle w,x\rangle\neq 0\\ \star&\text{otherwise}\end{cases}.$
Note that for every $\gamma>0$, $\mathcal{H}^{d}_{\gamma}$ is a subclass of $\mathcal{H}^{d}$.
Since $\mathcal{H}^{2}$ has infinite threshold dimension [AHHM22, Definition 28], by the celebrated result of [ABL+22], it is not globally stable and therefore $\operatorname{LR}(\mathcal{H}^{2})=\infty$. Consequently, the aforementioned result of [CMY23] implies the following interesting gap:
$\operatorname{LR}(\mathcal{H}^{2})=\infty\qquad\text{while}\qquad\widetilde{\operatorname{LR}}(\mathcal{H}^{2})=2.$
The authors of [CMY23] ask whether a similar bound holds for $\widetilde{\operatorname{LR}}(\mathcal{H}^{3})$. In the following theorem, we answer their question by extending their result to all dimensions.
Theorem 1.11.
For every dimension $d>1$, we have $\widetilde{\operatorname{LR}}(\mathcal{H}^{d})=d$.
Proof.
The upper bound is an easy consequence of the upper bound of our main theorem (Theorem 1.6). Indeed, any finite set of points $S\subset\mathbb{S}^{d-1}$ and hypotheses $H\subset\mathcal{H}^{d}|_{S}$ defined by unit vectors $W\subset\mathbb{S}^{d-1}$ is a sub-concept class of $\mathcal{H}^{d}_{\gamma}$, where $\gamma=\min_{x\in S,w\in W}|\langle w,x\rangle|$. The claim now follows from our upper bound from Theorem 1.6.
For the lower bound, we use a result of Chase, Moran, and Yehudayoff [CMY23, Theorem 3] stating that for every concept class $\mathcal{H}$,
$\operatorname{LR}(\mathcal{H})\geq\operatorname{VCdim}(\mathcal{H}).$
The result now follows as $\operatorname{VCdim}(\mathcal{H}^{d})=d$, and therefore, $\mathcal{H}^{d}$ has a finite subclass of VC dimension $d$. ∎
The sign-rank of a sign matrix $A$, denoted by ${\mathbf{rk}}_{\pm}(A)$, is the smallest rank of a real matrix $B_{\mathcal{X}\times\mathcal{Y}}$ such that the entries of $B$ are nonzero and have the same signs as their corresponding entries in $A$. The definition of sign-rank naturally extends to partial matrices, where for invalid entries of $A$, the corresponding entry in $B$ could be any real number.
Geometrically, sign-rank is the smallest dimension in which the matrix is realized as points and homogeneous half-spaces. We will state the definition of sign-rank in this terminology for partial classes.
Definition 1.12 (Sign-rank).
The sign-rank of a partial class {±,}$\mathcal{H}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$, denoted by ${\mathbf{rk}}_{\pm}(\mathcal{H})$, is the smallest $d$ such that there exist vectors $u_{h},v_{x}\in\mathbb{R}^{d}$ for all pairs $h\in\mathcal{H},x\in\mathcal{X}$ such that $h(x)=\operatorname{sgn}({\left\langle u_{h},v_{x}\right\rangle})$ whenever $h(x)\neq\star$.
Theorem 1.11 and the VC lower bound of [CMY23, Theorem 3] immediately imply the following general bounds on $\widetilde{\operatorname{LR}}(\mathcal{H})$.
Corollary 1.13.
For every partial class {±,}$\mathcal{H}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$, we have
$\operatorname{VCdim}(\mathcal{H})\leq\widetilde{\operatorname{LR}}(\mathcal{H})\leq\operatorname{rank_{\pm}}(\mathcal{H}).$
Gap Hamming Distance.
The discrete analogue of large-margin half-spaces is the well-studied Gap Hamming Distance (GHD) problem, a central problem in communication complexity. For $n\in\mathbb{N}$ and $\gamma\in(0,1)$, the $n$-bit $\operatorname{GHD}_{\gamma}$ problem, denoted $\operatorname{GHD}_{\gamma}^{n}$, is the partial function on inputs $x,y\in\{\pm 1\}^{n}$ defined by
$\operatorname{GHD}_{\gamma}^{n}(x,y)\coloneqq\begin{cases}\operatorname{sgn}({\left\langle x,y\right\rangle})&\text{if $|{\left\langle x,y\right\rangle}|\geq\gamma n$}\\ \star&\text{otherwise}\end{cases}.$
For fixed $\gamma\in(0,1)$, the public-coin randomized communication complexity of $\operatorname{GHD}_{\gamma}$ is at most a fixed constant depending only on $\gamma$. This places $\operatorname{GHD}_{\gamma}$ in the complexity class $\sf{BPP}_{0}$, which consists of communication problems with constant-cost public-coin randomized protocols. Moreover, $\operatorname{GHD}_{\gamma}$ exhibits certain properties that separate $\sf{BPP}_{0}$ from some complexity classes (e.g., $\sf{UPP}_{0}$) and measures [HHH23, CLV19, HHM23, FGHH25, Son14].
A key caveat of these separations is that $\operatorname{GHD}_{\gamma}$ is a partial function, and whether it extends to any family of total functions with similar properties remained an open problem before this work. One suggested approach [FGHH25] to resolving these questions is to find a disambiguation of $\operatorname{GHD}_{\gamma}$ to a total function with constant public-coin randomized communication complexity. In Corollary 1.15 below, we prove this is impossible: every disambiguation of $\operatorname{GHD}_{\gamma}^{n}$ has public-coin randomized communication complexity at least $\Omega(\log\log n)$.
First, we show that every such disambiguation has an unbounded Littlestone dimension.
Theorem 1.14.
Let $\gamma\in(0,1)$ be a margin parameter. Every family of disambiguations $\{M_{n}\}_{n=1}^{\infty}$ of the Gap Hamming Distance matrices $\{\mathrm{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$ satisfies
$\operatorname{Ldim}(M_{n})=\Omega(\sqrt{\log n}).$
We obtain the following corollary by combining Theorem 1.14 with some known results.
Corollary 1.15.
Let $\gamma\in(0,1)$ be a margin parameter. Every family of disambiguation of the Gap Hamming Distance matrices $\{\mathrm{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$ has public-coin randomized communication complexity $\Omega(\log\log n)$.
The bound follows by combining Theorem 1.14 with the known relationship between Littlestone dimension, margin, distributional discrepancy, and public-coin randomized communication complexity. See Section 4 for the proof.
1.2 Concluding remarks and open problems
While Theorem 1.4 provides an elegant and comprehensive characterization of various notions of replicability and privacy for total concepts through the combinatorial framework of the Littlestone dimension, the landscape of replicability for partial classes, as demonstrated by the results of this paper, is more intricate and less understood.
The “DP-learnability to Shared-randomness Replicability” reduction from [BGH+23] extends to the partial setting. Moreover, [FHM+24] recently showed that for partial classes, DP-learnability implies a finite Littlestone dimension.
Theorem 1.16 ([BGH+23, FHM+24]).
Let {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ be a partial concept class.
  • If $\mathcal{H}$ is DP-learnable, then $\mathcal{H}$ is shared-randomness replicable.
  • If $\mathcal{H}$ is DP-learnable, then $\operatorname{Ldim}(\mathcal{H})<\infty$.
On the other hand, our main theorem shows that for partial concepts, list replicability does not follow from bounded Littlestone dimension, shared-randomness replicability, DP-learnability, or even pure DP-learnability. It is also straightforward to show that, even in the partial setting, list replicability implies shared-randomness replicability.
Proposition 1.17 ([KKMV23, Lemma 8]).
If {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is list replicable (equiv. globally stable), then $\mathcal{H}$ is shared-randomness replicable.
To our knowledge, no further relationships among the four notions of DP-learnability, shared-randomness replicability, Littlestone dimension, and list replicability are known for partial functions.
Open problems.
Finally, we list some open problems for future research that naturally arise from our work.
  1. 1.
    Are DP-learnability, shared-randomness replicability, and finite Littlestone dimension equivalent for partial functions? If not, what are the precise relationships between them?
  2. 2.
    Is there a combinatorial notion of dimension that characterizes list replicability?
  3. 3.
    How tight are the inequalities in Corollary 1.13, namely,
    $\operatorname{VCdim}(\mathcal{H})\leq\widetilde{\operatorname{LR}}(\mathcal{H})\leq\operatorname{rank_{\pm}}(\mathcal{H})?$
  4. 4.
    Is the answer to 1.7 negative for VC dimension?
    In an article posted on arXiv just a few days ago, Chornomaz, Moran, and Waknine [CMW25] explored this problem using a topological approach, but the question remains open.
2 Preliminaries
Much of the notation and definitions were already outlined in the introduction. In this section, we review a few key concepts from learning theory, specifically PAC learnability, the VC and Littlestone dimensions, and differential privacy.
In some of our proofs, we will apply Hoeffding’s classic concentration inequality, which we state below for reference.
Theorem 2.1 (Hoeffding’s inequality).
Let $c\in\mathbb{R}$ and let $\bm{x}_{1},\ldots,\bm{x}_{n}$ be independent random variables with $\bm{x}_{i}\in[-c,c]$ and $\mathbb{E}[\bm{x}_{i}]=0$. Then, for any $t>0$,
$\Pr\left[\left|\sum_{i=1}^{n}\bm{x}_{i}\right|\geq t\right]\leq 2e^{-\frac{t^{2}}{2c}}.$
We define the population loss of a partial concept $h\in\left\{\pm 1,\star\right\}^{\mathcal{X}}$ with respect to a distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$ as
$\mathcal{L}_{\mathcal{D}}(h)\coloneqq\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}[h(\bm{x})\neq\bm{y}].$
Under this definition, predictions with $h(x)=\star$ always count as mistakes. The population loss of a class $\mathcal{H}$ is given by the smallest population loss achievable within $\mathcal{H}$:
$\mathcal{L}_{\mathcal{D}}(\mathcal{H})\coloneqq\inf_{h\in\mathcal{H}}\mathcal{L}_{\mathcal{D}}(h).$
The following simple lemma from [AHHM22] establishes the connection between realizability and having zero population loss.
Lemma 2.2 ([AHHM22]).
Let {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ be a partial concept class, and let $\mathcal{D}$ be a distribution on $\mathcal{X}\times\left\{\pm 1\right\}$. If $\mathcal{L}_{\mathcal{D}}(\mathcal{H})=0$, then $\mathcal{D}$ is realizable by $\mathcal{H}$. Conversely, if $\mathcal{D}$ is realizable and has finite or countable support, then $\mathcal{L}_{\mathcal{D}}(\mathcal{H})=0$.
PAC learning of partial classes:
Given a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$, the goal of PAC (Probably Approximately Correct) learning is for the learning rule to produce, with high probability, a hypothesis whose population loss is close to the best achievable loss within $\mathcal{H}$.
Definition 2.3 (PAC learning of partial concept classes).
A class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is PAC learnable if there is a learning rule $\bm{\mathcal{A}}$ and a function $n(\epsilon,\delta)$ such that for every $\epsilon,\delta>0$ and every realizable distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm\right\}$, we have
$\Pr_{\bm{S}\sim\mathcal{D}^{n}}[\mathcal{L}_{\mathcal{D}}(\bm{\mathcal{A}}(\bm{S}))\leq\epsilon]\geq 1-\delta\ \text{ where }n=n(\epsilon,\delta).$
2.1 VC and Littlestone dimensions
We now extend the definition of the VC dimension to partial concept classes.
Definition 2.4 (Shattered set).
A finite set of points $C=\left\{x_{1},\ldots,x_{n}\right\}\subseteq\mathcal{X}$ is shattered by a partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ if, for every pattern $y\in\left\{\pm 1\right\}^{n}$, there exists $h\in\mathcal{H}$ such that $h(x_{i})=y_{i}$ for all $i\in[n]$.
Definition 2.5 (VC dimension).
The VC dimension of a partial concept class $\mathcal{H}$, denoted $\operatorname{VCdim}(\mathcal{H})$, is the largest integer $d$ for which there exists a size-$d$ subset of $\mathcal{X}$ that is shattered by $\mathcal{H}$. If no such maximum $d$ exists, we define $\operatorname{VCdim}(\mathcal{H})=\infty$.
Viewed as a matrix, the VC dimension of $\mathcal{H}$ is the largest $d$ such that the associated partial matrix $M_{\mathcal{X}\times\mathcal{H}}$ contains a $d\times 2^{d}$ submatrix with ±$\pm 1$ entries, where the columns enumerate all $d$-bit ±$\pm 1$ patterns.
The fundamental theorem of PAC learning asserts that a total concept class is PAC learnable if and only if its VC dimension is finite. This result was extended to partial concept classes by [AHHM22] using a completely different proof.
Theorem 2.6 ([AHHM22]).
The following statements are equivalent for any partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$.
  • $\operatorname{VCdim}(\mathcal{H})<\infty$;
  • $\mathcal{H}$ is PAC learnable;
Littlestone dimension:
In his influential work, Littlestone [Lit88] provided a combinatorial characterization of concept classes that are online learnable. To this end, he introduced a measure that extends the VC dimension by considering the shattering of decision trees rather than sets.
A mistake tree of depth $d$ over a domain $\mathcal{X}$ is a complete binary tree of depth $d$ with the following properties:
  • Each internal node is labeled by an element $x\in\mathcal{X}$.
  • Each edge is labeled by a binary value $b\in\left\{\pm 1\right\}$, where $b=-1$ indicates a left child and $b=1$ indicates a right child.
Every root-to-leaf path in the tree is described by a sequence
$(x_{1},b_{1}),\ldots,(x_{d},b_{d}),$
where $x_{i}\in\mathcal{X}$ is the label of the $i$th internal node along the path, and $b_{i}$ specifies whether the path moves to the left or right child at each level.
A concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ shatters a mistake tree if, for every root-to-leaf path
$(x_{1},b_{1}),\ldots,(x_{d},b_{d}),$
there exists a hypothesis $h\in\mathcal{H}$ such that $h(x_{i})=b_{i}$ for all $i\in[d]$.
Definition 2.7 (Littlestone Dimension).
The Littlestone dimension of a concept class $\mathcal{H}$, denoted $\operatorname{Ldim}(\mathcal{H})$, is the largest integer $d$ for which there exists a mistake tree of depth $d$ that is shattered by $\mathcal{H}$. If no such maximum $d$ exists, we define $\operatorname{Ldim}(\mathcal{H})=\infty$.
It always holds that $\operatorname{VCdim}(\mathcal{H})\leq\operatorname{Ldim}(\mathcal{H})$, since any set $C=\left\{x_{1},\ldots,x_{d}\right\}$ shattered by $\mathcal{H}$ gives rise to a mistake tree of depth $d$, where all nodes at level $i$ are labeled with $x_{i}$. This tree is shattered by $\mathcal{H}$.
Littlestone proved that a total concept class $\mathcal{H}$ is online learnable if and only if $\operatorname{Ldim}(\mathcal{H})<\infty$. This result was later extended to partial concept classes by [AHHM22].
2.2 Differentially Private Learnability
The widely adopted approach for ensuring privacy in machine learning is the differential privacy (DP) framework, introduced in [DMNS06]. Informally, differential privacy in learning means that no single labeled example in the input dataset significantly impacts the learner’s output hypothesis. In other words, the output distribution of a differentially private randomized learning algorithm remains nearly unchanged if a single data point is modified.
Differential privacy is quantified with two parameters $\epsilon,\delta>0$. We say that two probability distributions $p$ and $q$ are $(\epsilon,\delta)$-indistinguishable, if for every event $E$, we have
$p(E)\leq e^{\epsilon}q(E)+\delta\ \text{ and }\ q(E)\leq e^{\epsilon}p(E)+\delta.$
Two random variables are $(\epsilon,\delta)$-indistinguishable if their distributions satisfy this condition.
Definition 2.8 (Differential Privacy).
Given $\epsilon,\delta>0$, a randomized learning rule
$\bm{\mathcal{A}}:(\mathcal{X}\times\left\{\pm 1\right\})^{n}\to\left\{\pm\right\}^{\mathcal{X}}$
is $(\epsilon,\delta)$-differentially-private if for every two samples $S,S^{\prime}\in(\mathcal{X}\times\left\{\pm\right\})^{n}$ differing on a single example, the random variables $\bm{\mathcal{A}}(S)$ and $\bm{\mathcal{A}}(S^{\prime})$ are $(\epsilon,\delta)$-indistinguishable.
We emphasize that $(\epsilon,\delta)$-indistinguishability must hold for every such pair of samples, regardless of whether they are drawn from a (realizable) distribution.
The special case where $\delta=0$ is known as pure differential privacy, while the more general case where $\delta>0$ is referred to as approximate differential privacy.
In approximate differential privacy, the parameters $\epsilon$ and $\delta$ are typically set as follows: $\epsilon$ is taken to be a small fixed constant (e.g., 0.1), while $\delta$ is a negligible function, $\delta=n^{\omega(1)}$, where $n$ is the sample size. Definitions 2.9 and 2.10 formally define differentially private learnability under the assumption that the privacy parameter $\epsilon$ is fixed at $\epsilon=0.1$.
Definition 2.9 (Approximate Differentially Private Learnability).
We say that a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is approximate differentially private learnable (DP-learnable) if there is a learning rule $\bm{\mathcal{A}}:(\mathcal{X}\times\left\{\pm 1\right\})^{*}\to\left\{\pm 1\right\}^{\mathcal{X}}$ with sample complexity $n(\epsilon,\delta,\delta^{\prime})$ such that the following holds.
  • For every $\epsilon,\delta,\delta^{\prime}>0$, the class $\mathcal{H}$ is $(\epsilon,\delta)$-PAC learnable by $\bm{\mathcal{A}}$ using $n(\epsilon,\delta,\delta^{\prime})$ samples.
  • For every $\epsilon,\delta,\delta^{\prime}>0$, the learning rule $\bm{\mathcal{A}}$ applied to samples of size $n(\epsilon,\delta,\delta^{\prime})$ is $(0.1,\delta^{\prime})$-differentially private.
  • For every fixed $\epsilon,\delta>0$, we have
    $\lim_{\delta^{\prime}\to 0}\frac{\log(1/\delta^{\prime})}{\log n(\epsilon,\delta,\delta^{\prime})}=\infty.$
Definition 2.10 (Pure Differentially Private Learnability).
We say that a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is pure differentially private learnable (pure DP-learnable) if $\mathcal{H}$ is PAC learnable by a $(0.1,0)$-differentially private learning rule.
3 Proof of the Main Theorem
3.1 The Lower Bound
We prove the lower bound via a topological argument that utilizes the following local version of the Borsuk-Ulam theorem proved in [CCMY24].
Theorem 3.1 (Local Borsuk-Ulam [CCMY24]).
Let $d\geq 2$ be an integer. If $\mathcal{F}$ is a finite antipodal-free open cover of the sphere $\mathbb{S}^{d-1}$, then there exists some $w\in\mathbb{S}^{d-1}$ contained in at least $\lceil\frac{d}{2}+1\rceil$ member sets of $\mathcal{F}$.
Fix any margin $\gamma\in(0,1)$, dimension $d\geq 2$ and $\epsilon\in(0,1/2)$, and suppose that $\bm{\mathcal{A}}$ is an $(\epsilon,L)$-list replicable learning rule for $\mathcal{H}_{\gamma}^{d}$. We prove that $L\geq\frac{d}{2}+1$.
By the definition of list replicability, for any $\delta>0$, there is an integer $n$ so that for any realizable distribution $\mathcal{D}$, there exists a list of hypotheses $\{h_{1},\dots,h_{L}\}$ with
$\operatorname{Pr}_{\mathbf{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\mathbf{S})\in\{h_{1},\dots,h_{L}\}]\geq 1-\delta\text{ and }\mathcal{L}_{\mathcal{D}}(h_{i})\leq\epsilon\text{ for all }i\in[L].$
Now pick any $\alpha>0$ and $\epsilon^{\prime}\in(\epsilon,1/2)$. By taking $\delta$ sufficiently small, for any distribution $\mathcal{D}$, we can choose a hypothesis $h_{\mathcal{D}}\in\{h_{1},\dots,h_{L}\}$ such that
$\Pr_{\mathbf{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{\mathcal{D}}]>\frac{1}{L+\alpha}\text{ and }\mathcal{L}_{\mathcal{D}}(h_{\mathcal{D}})<\epsilon^{\prime}.$ (3)
We will focus on a certain collection of realizable distributions $\mathcal{D}$ on $\mathbb{S}^{d-1}\times\{\pm 1\}$. For any $w\in\mathbb{S}^{d-1}$, take $\mathcal{D}_{w}$ to be the uniform distribution on the set $\{(x,h_{w}(x))\ |\ x\in{\mathrm{supp}}(h_{w})\}$. These distributions are, by definition, realizable. Hence, for each $\mathcal{D}_{w}$, we can choose some particular hypothesis $h_{\mathcal{D}_{w}}$ that satisfies the conditions in Eq. 3. Collect these hypotheses in a set $T$, that is to say
$T\coloneqq\{h_{\mathcal{D}_{w}}\ |\ w\in\mathbb{S}^{d-1}\}.$
For each $h\in T$, define the set $C_{h}\subset\mathbb{S}^{d-1}$ as
$C_{h}\coloneqq\left\{w\in\mathbb{S}^{d-1}\ |\ \Pr_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h]>\frac{1}{L+\alpha}\text{ and }\mathcal{L}_{\mathcal{D}_{w}}(h)<\epsilon^{\prime}\right\}.$
Claim 3.2.
The family $\{C_{h}\}_{h\in T}$ forms an antipodal-free open cover of $\mathbb{S}^{d-1}$.
Proof.
The fact that any set $C_{h}$ is antipodal-free follows from the accuracy constraint $\mathcal{L}_{\mathcal{D}_{w}}(h)<\epsilon^{\prime}$. Indeed, for any $w\in\mathbb{S}^{d-1}$, the concepts $h_{w}$ and $h_{-w}$ have identical support, on which they disagree at every point. Thus the population loss of any hypothesis $h$ satisfies the equation
$\mathcal{L}_{\mathcal{D}_{w}}(h)+\mathcal{L}_{\mathcal{D}_{-w}}(h)=1.$
For any $w\in C_{h}$, we have that $\mathcal{L}_{\mathcal{D}_{w}}(h)<\epsilon^{\prime}<1/2$, whereby $w$ and $-w$ cannot both be in $C_{h}$.
Next, each set $C_{h}$ is open because both $\Pr_{\bm{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\bm{S})=h]$ and $\mathcal{L}_{\mathcal{D}_{w}}(h)$ are continuous in $w$. Lastly, the family $\{C_{h}\}_{h\in T}$ covers $\mathbb{S}^{d-1}$ because, for any $w\in\mathbb{S}^{d-1}$, the set $C_{h_{\mathcal{D}_{w}}}$ contains $w$ by construction. ∎
Now note that the antipodal-free open cover $\{C_{h}\}_{h\in T}$ admits a finite subcover by the compactness of the unit sphere. Applying Theorem 3.1 to such a finite subcover guarantees that some $w\in\mathbb{S}^{d-1}$ is contained in at least $t\coloneqq\lceil\frac{d}{2}+1\rceil$ sets $C_{h_{1}},C_{h_{2}},\dots,C_{h_{t}}$. Unpacking definitions reveals that the distribution $\mathcal{D}_{w}$ has the property
$\Pr_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]>\frac{1}{L+\alpha}$
for $t$ distinct hypotheses $h_{i}\in T$. Because these $h_{i}$ are distinct, the events $[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]$ are disjoint, and therefore
$1\geq\operatorname{Pr}_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}\bigcup_{i=1}^{t}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]=\sum_{i=1}^{t}\operatorname{Pr}_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]>\frac{t}{L+\alpha}.$
It follows that $L+\alpha>t=\lceil\frac{d}{2}+1\rceil$ for any $\alpha>0$, which implies the desired lower bound $L\geq\lceil\frac{d}{2}+1\rceil$.
3.2 The Upper Bound
To prove the upper bound, we design a list replicable learning algorithm $\bm{\mathcal{A}}$ that learns $\mathcal{H}_{\gamma}^{d}$ with list size $d$ independent of $\epsilon>0$. Given $w\in\mathbb{S}^{d-1}$, let $\overline{h}_{w}:\mathbb{S}^{d-1}\to\{\pm 1\}$ denote the total concept class corresponding to the closed half-space defined by $w$.
$\overline{h}_{w}(x)\coloneqq\begin{cases}1&\text{if }\langle w,x\rangle\geq 0\\ -1&\text{if }\langle w,x\rangle<0\end{cases}.$
Fundamentally, as in [KKL+24], we estimate a large-margin linear separator using the average of many runs of an SVM maximum margin separator. Then, we use a rounding scheme based on a uniform triangulation of the $\ell_{1}$ sphere, with the guarantee that with high probability, our learning rule will choose one of at most $d$ separators.
Consider a training sample $(x_{1},y_{1}),\dots,(x_{n},y_{n})\in\mathbb{R}^{d}\times\{\pm 1\}$. The (homogeneous) hard-SVM is an optimization problem that returns a homogeneous half-space that classifies the training sample correctly while maximizing the margin $\gamma$. More formally, it is the following optimization problem over the variables $\gamma\in\mathbb{R}$ and $w\in\mathbb{S}^{d-1}$:
$\begin{array}[]{l rll}\max&\gamma&&\\ \text{s.t. }&y_{i}{\left\langle x_{i},w\right\rangle}&\geq\gamma&\text{for }i=1,\ldots,n\\ &\gamma&\geq 0&\\ &w&\in\mathbb{S}^{d-1}&\\ \end{array}$
One can use semi-definite programming to solve this optimization problem efficiently—to check whether it is feasible and, if so, to find the maximizing $w$.
Definition 3.3 ($\gamma$-Separator).
Let $S\subseteq\mathbb{S}^{d-1}\times\left\{\pm 1\right\}$. We call $w\in\mathbb{S}^{d-1}$ a $\gamma$-separator for $S$ if
$y{\left\langle x,w\right\rangle}\geq\gamma\ \ \text{ for all }(x,y)\in S.$
Furthermore, for any distribution $\mathcal{D}$ over $\mathbb{S}^{d-1}\times\{\pm 1\}$, we call $w$ a $(\gamma,\epsilon)$-separator for $\mathcal{D}$ if
$\Pr_{(x,y)\sim\mathcal{D}}[y\langle x,w\rangle<\gamma]\leq\epsilon.$
When learning $\mathcal{H}_{\gamma}^{d}$ in the realizable setting, for any sample set $S$ drawn from a realizable distribution $\mathcal{D}$, there is some $w$ that $\gamma$-separates $S$. Therefore, the hard-SVM will be feasible and return a vector $w$ that $\gamma$-separates $S$.
The following theorem, due to [STBWA98], says that if we take a sufficiently large sample $\bm{S}$ and compute a good separator $w$ for it using hard-SVM, then with high probability, $w$ will also be a good separator for $\mathcal{D}$.
Theorem 3.4 (SVM generalization bound [STBWA98, Theorem 3.5]).
For all $\epsilon,\delta>0$, there exists $n\coloneqq n(\epsilon,\delta)$ such that the following holds. Let $\mathcal{D}$ be any distribution over $\mathbb{S}^{d-1}\times\left\{\pm 1\right\}$.
$\Pr_{\bm{S}\sim\mathcal{D}^{n}}\left[\text{Every $w\in\mathbb{S}^{d-1}$ that $\gamma$-separates $\bm{S}$ also $\left(\frac{\gamma}{2},\epsilon\right)$-separates $\mathcal{D}$}\right]\geq 1-\delta.$
Remark 3.5.
To prove Theorem 3.4, one can apply [STBWA98, Theorem 3.5] to show that, with probability at least $1-\delta$, both $h_{1}(x)\coloneqq\operatorname{sgn}({\left\langle x,w\right\rangle}+\frac{\gamma}{2})$ and $h_{2}(x)\coloneqq\operatorname{sgn}({\left\langle x,w\right\rangle}-\frac{\gamma}{2})$ have loss at most $\frac{\epsilon}{2}$, in which case $w$ is a $\left(\frac{\gamma}{2},\epsilon\right)$-separator for $\mathcal{D}$.
Regarding optimal bounds on $n(\epsilon,\delta)$ in Theorem 3.4, we refer the reader to [GKL20, KKL+24].
We will also use the following simple concentration result for sums of i.i.d. random vectors to show that the outputs of multiple runs of hard-SVM on independent samples are typically concentrated around their mean.
Lemma 3.6.
Let $\bm{x}_{1},\dots,\bm{x}_{k}\in\mathbb{R}^{d}$ be i.i.d random variables with mean $\mu$ and $\|\bm{x}_{i}-\mu\|_{\infty}\leq C$. Let $\bm{Z}=\frac{1}{k}\sum_{i=1}^{k}\bm{x}_{i}$. For all $t>0$,
$\Pr[\|\bm{Z}-\mu\|_{1}\geq t]\leq 2de^{\frac{-kt^{2}}{2d^{2}C^{2}}}.$
Proof.
We apply Hoeffding’s inequality to each coordinate and take the union bound. By Hoeffding’s inequality, for every $j\in[d]$, we have
$\Pr\left[|\bm{Z}_{j}-\mu_{j}|\geq\frac{t}{d}\right]\leq 2e^{\frac{-kt^{2}}{2d^{2}C^{2}}}.$
Therefore, by the union bound,
$\Pr\left[\|\bm{Z}-\mu\|_{1}\geq t\right]\leq 2de^{\frac{-kt^{2}}{2d^{2}C^{2}}}.\qed$
We will use a rounding scheme that ensures any small neighbourhood on $\mathbb{S}^{d-1}$ is rounded to at most $d$ points.
Lemma 3.7.
For every $\alpha>0$, there is a $\beta(d)>0$ and a rounding scheme
$\mathrm{round}_{\alpha}:\mathbb{S}^{d-1}\to\mathbb{S}^{d-1}$
such that for all $x\in\mathbb{S}^{d-1}$,
  1. 1.
    $\|\mathrm{round}_{\alpha}(x)-x\|_{2}<\alpha$, and
  2. 2.
    The set $R_{x}\coloneqq\left\{\mathrm{round}_{\alpha}(y)\ |\ y\in\mathbb{S}^{d-1}\text{and }\|x-y\|_{2}\leq\beta\right\}$ has size at most $d$.
Proof.
Consider any $\frac{\alpha}{2}$-net $T$ of points in general position on $\mathbb{S}^{d-1}$ and define the rounding function as
$\mathrm{round}_{\alpha}(x)\coloneqq\arg\min_{y\in T}\|x-y\|_{2}.$
Property 1 follows directly from the definition of $\mathrm{round}_{\alpha}$, so it remains to prove 2.
If $|T|\leq d$, both conditions are satisfied. Thus, assume $|T|>d$. We will use the fact that for any set of $d+1$ distinct points $x_{1},\ldots,x_{d+1}\in T$, the origin is the only point equidistant from all of them. To see this, suppose there exists a point $y\in\mathbb{R}^{d}$ that is equidistant from each $x_{i}$, meaning there exists some $r$ such that
$r^{2}=\|x_{i}-y\|_{2}^{2}=1+\|y\|_{2}^{2}-2{\left\langle x_{i},y\right\rangle}.$
Consequently, $y$ is orthogonal to the linearly independent vectors $x_{1}-x_{2}$, …, $x_{1}-x_{d+1}$, and thus $y=\vec{0}$.
Define the map $\phi:\mathbb{S}^{d-1}\to\mathbb{R}_{\geq 0}$ as
$\phi(x)\coloneqq\tau(x)-\min_{y\in T}\|x-y\|_{2},$
where $\tau(x)$ denotes the distance from $x$ to a $(d+1)$-th closest point in $T$. Since no point in $\mathbb{S}^{d-1}$ can be equidistant to more than $d$ points in $T$, we have $\phi(x)>0$ for all $x$. And since $\phi$ is continuous and $\mathbb{S}^{d-1}$ is compact, we have
$\beta^{\prime}\coloneqq\min_{x}\phi(x)>0.$
Taking $\beta\coloneqq\beta^{\prime}/3$ completes the proof. ∎
Upper bound of Theorem 1.6.
We need to show that for any margin $\gamma\in(0,1)$, accuracy parameter $\epsilon\in(0,1/2)$, and dimension $d\geq 1$, we have $\operatorname{LR}_{\epsilon}(\mathcal{H}_{\gamma}^{d})\leq d$.
We will construct a list-replicable learner that always outputs a hypothesis of the form $\overline{h}_{w}$ for some $w\in\mathbb{S}^{d-1}$.
Let $k=k(d,\gamma)$ and $n_{0}=n_{0}(\epsilon,\delta,k)$ be integers yet to be determined. Consider the following learning rule $\bm{\mathcal{A}}$ that uses the rounding scheme of Lemma 3.7.
Algorithm 1 The learning rule $\bm{\mathcal{A}}$
1:  for $i\leftarrow 1$ to $k$ do
2:     Sample $\mathbf{S}_{i}\sim\mathcal{D}^{n_{0}}$.
3:     Let $\bm{w}_{i}\leftarrow\text{hard-SVM}(\mathbf{S}_{i})$.
4:  end for
5:  Let $\bm{w}\leftarrow\frac{1}{k}\sum_{i=1}^{k}\bm{w}_{i}$ and $\bm{z}\leftarrow\frac{\bm{w}}{\|\bm{w}\|_{2}}$.
6:  Let $\tilde{\bm{z}}\leftarrow\mathrm{round}_{\gamma/2}(\bm{z})$.
7:  output the hypothesis $\overline{h}_{\tilde{\bm{z}}}$.
We first show that the learning rule $\bm{\mathcal{A}}$ presented in Algorithm 1 is a PAC learner.
Claim 3.8.
Let $\bm{\mathcal{A}}$ and $\bm{w}$ be as in Algorithm 1. For every $\epsilon,\delta\in(0,1)$ and $k\in\mathbb{N}$, there exists $n_{0}\coloneqq n_{0}(\epsilon,\delta,k)\in\mathbb{N}$ such that for every distribution $\mathcal{D}$ realizable by $\mathcal{H}_{\gamma}^{d}$, we have
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{w}\|_{2}<\frac{\gamma}{2}\right]\leq\frac{\delta}{4}$ (4)
and
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}[\mathcal{L}_{\mathcal{D}}(\bm{\mathcal{A}}(\mathbf{S}))\geq\epsilon]\leq\frac{\delta}{4}.$ (5)
Proof.
Let $\bm{w}_{1},\dots,\bm{w}_{k}$ be as in Algorithm 1. Since $\mathcal{D}$ is realizable by $\mathcal{H}_{\gamma}^{d}$, for every $i$, $\bm{w}_{i}$ is a $\gamma$-separator for $\mathbf{S}_{i}$. Therefore, by Theorem 3.4, if $n_{0}(\epsilon,\delta,k)$ is sufficiently large,
$\Pr_{\begin{subarray}{c}\mathbf{S}_{i}\sim\mathcal{D}^{n_{0}}\end{subarray}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\frac{\epsilon}{k}\right]\geq 1-\frac{\delta}{4k}.$
Thus, by the union bound,
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\frac{\epsilon}{k}\text{ for all $i\in[k]$}\right]\geq 1-\frac{\delta}{4},$ (6)
and applying the union bound again,
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\min_{i\in[k]}\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\epsilon\right]\geq 1-\frac{\delta}{4}.$ (7)
Finally, if $(x,y)\in\mathbb{S}^{d-1}\times\left\{\pm 1\right\}$ satisfies $y{\left\langle x,\bm{w}_{i}\right\rangle}\geq\gamma/2$ for all $i\in[k]$, then noting that $\|\bm{w}\|\leq 1$, we have
$y{\left\langle x,\bm{z}\right\rangle}\geq y{\left\langle x,\bm{w}\right\rangle}=y{\left\langle x,\frac{1}{k}\sum_{i=1}^{k}\bm{w}_{i}\right\rangle}\geq\gamma/2.$ (8)
Thus, from Equation 7, we have
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{w}\|_{2}<\frac{\gamma}{2}\right]\leq\frac{\delta}{4}$
and
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\bm{y}{\left\langle\bm{x},\bm{z}\right\rangle}\geq\frac{\gamma}{2}\right]\geq 1-\epsilon\right]\geq 1-\frac{\delta}{4}.$ (9)
By applying Lemma 3.7 with $\alpha\coloneqq\gamma/2$, after rounding $\bm{z}$ to $\tilde{\bm{z}}$, we have $\|\tilde{\bm{z}}-\bm{z}\|_{2}<\frac{\gamma}{2}$. Thus if $(x,y)\in\mathbb{S}^{d-1}\times\{\pm 1\}$ satisfy $y{\left\langle x,\bm{z}\right\rangle}\geq\gamma/2$, then
$y{\left\langle x,\tilde{\bm{z}}\right\rangle}=y{\left\langle x,\bm{z}\right\rangle}+y{\left\langle x,\tilde{\bm{z}}-\bm{z}\right\rangle}\geq\frac{\gamma}{2}-\|\tilde{\bm{z}}-\bm{z}\|_{2}>0,$
namely $\overline{h}_{\tilde{\bm{z}}}(x)=y$. Thus,
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\mathcal{L}_{\mathcal{D}}(\overline{h}_{\tilde{\bm{z}}})\leq\epsilon\right]\geq 1-\frac{\delta}{4},$
which completes the proof of the claim. ∎
We now complete the proof by addressing list replicability. Let $\beta$ be as in Lemma 3.7. Applying Lemma 3.6, since $\bm{w}$ is the average of $k$ i.i.d. random variables in $\mathbb{S}^{d-1}$, there exists $k=k(\gamma,d)\in\mathbb{N}$ such that
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{w}-\operatorname{\mathbb{E}}[\bm{w}]\|_{2}\geq\frac{\gamma\beta}{2}\right]\leq\frac{\delta}{4}.$
Since $\bm{z}=\frac{\bm{w}}{\|\bm{w}\|_{2}}$, by applying the union bound to Eq. 4 and the above inequality, we have
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{z}-\operatorname{\mathbb{E}}[\bm{z}]\|_{2}\geq\beta\right]\leq\frac{\delta}{2}.$
Consequently, by Lemma 3.7, with probability at least $21-\delta/$ 2, the rounding scheme ($\mathrm{round}_{\frac{\gamma}{2}}$) outputs one of at most $d$ hypotheses. Applying a union bound with 3.8 completes the proof of the upper bound of Theorem 1.6. ∎
4 Disambiguations of Gap Hamming Distance
We first prove that every disambiguation of the Gap Hamming Distance matrix family has unbounded public-coin randomized communication complexity.
Proof of Corollary 1.15.
We combine Theorem 1.14 with the known relationship between Littlsetone dimension, margin, distributional discrepancy, and public-coin randomized communication complexity. Given a matrix $M\in\{\pm 1\}^{\mathcal{X}\times\mathcal{Y}}$, the margin of $M$ is defined
$\operatorname{m}(M)\coloneqq\max_{\begin{subarray}{c}d\in\mathbb{N},\\ u_{x},u_{y}\in\mathbb{S}^{d}\end{subarray}}\min_{(x,y)}M(x,y)\cdot{\left\langle u_{x},u_{y}\right\rangle}.$
In other words, $\operatorname{m}(M)$ is the largest $\gamma$ such that $M$ appears as a submatrix of $\mathcal{H}^{d}_{\gamma}$ for some $d$.
Let $\{M_{n}\}_{n=1}^{\infty}$ be a family of disambiguation of $\{\mathrm{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$. By Theorem 1.14, we know that $\operatorname{Ldim}(M_{n})=\Omega(\sqrt{\log n})$. It thus follows from LABEL:{Ldim_of_margin} that
$\operatorname{m}(M_{n})=O\left(\frac{1}{\sqrt[4]{\log n}}\right).$
Finally, invoking the equivalence of margin and discrepancy by [LS09] and the relation between discrepancy and randomized communication complexity by [CG88] (see also [HHP+22, Proposition 3.3 ]) shows that the public-coin randomized communication complexity of $M_{n}$ is
$\Omega(\log(\operatorname{m}(M_{n})^{-1}))=\Omega(\log\log n).\qed$
We next present the proof of Theorem 1.14. The key is to use an embedding of bounded margin half-spaces in dimension $d$ into the Boolean cube, which allows us to disambiguate $\mathcal{H}^{d}_{\gamma}$ using a disambiguation of the Gap Hamming Distance problem in dimension $O(d)$. The existence of such an embedding was proved in [HHM23], which we rephrase as follows.
Lemma 4.1 (Adapted from [HHM23, Lemma 3.2]).
Let $\gamma\in(0,1)$ and $n\in\mathbb{N}$. There exist $d=\Omega\left((1-\gamma)^{2}\cdot n/\log(1/(1-\gamma))\right)$, $\gamma^{\prime}\in(0,1)$ and a map $\xi\colon\mathbb{S}^{d-1}\to\left\{\pm 1\right\}^{n}$ such that for all $u,v\in\mathbb{S}^{d-1}$, we have
$\langle u,v\rangle\geq\gamma^{\prime}\implies\langle\xi(u),\xi(v)\rangle\geq\gamma n$
$\langle u,v\rangle\leq-\gamma^{\prime}\implies\langle\xi(u),\xi(v)\rangle\leq-\gamma n$
We are now ready to prove Theorem 1.14.
Proof of Theorem 1.14.
Fix $\gamma\in(0,1)$. Let $\{M_{n}\}_{n=1}^{\infty}$ be a family of total functions which disambiguates $\{\operatorname{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$, and let $d=d(n)$ and $\gamma^{\prime}$ be as provided by Lemma 4.1.
We will use this lemma along with the functions $\left\{M_{n}\right\}_{n=1}^{\infty}$ to disambiguate the family of partial concept classes $\{\mathcal{H}_{\gamma^{\prime}}^{d(n)}\}_{n=1}^{\infty}$. To this end, we disambiguate each partial concept $h_{w}\in\mathcal{H}_{\gamma^{\prime}}^{d}$ (defined in LABEL:{eq:def_large_margin}) to
$\overline{h}_{w}(x)\coloneqq M_{n}\big{(}\xi(w),\xi(x)\big{)}.$
Let us verify that $\overline{h}_{w}$ is, in fact, a disambiguation of $h_{w}$.
Suppose that $h_{w}(x)=1$. By definition, this occurs exactly when $\langle w,x\rangle\geq\gamma^{\prime}$. It follows from the properties of $\xi$ that
$\langle\xi(w),\xi(x)\rangle\geq\gamma n.$
Therefore, for such $w,x$, we have
$\overline{h}_{w}(x)=M_{n}(\xi(w),\xi(x))=\operatorname{GHD}_{\gamma}^{n}(\xi(w),\xi(x))=1=h_{w}(x).$
A similar argument shows that if $h_{w}(x)=-1$, then $\overline{h}_{w}(x)=-1$. We deduce that $\overline{h}_{w}$ indeed disambiguates $h_{w}$, and ¯$\overline{\mathcal{H}}^{d}_{\gamma^{\prime}}$ is a disambiguation of $\mathcal{H}^{d}_{\gamma^{\prime}}$.
Finally, note that by construction, any shattered mistake tree in ¯$\overline{\mathcal{H}}_{\gamma^{\prime}}^{d}$ corresponds to a shattered mistake tree of the same depth in $M_{n}$. Therefore, $\operatorname{Ldim}(\overline{\mathcal{H}}_{\gamma^{\prime}}^{d})\leq\operatorname{Ldim}(M_{n})$. This combined with Theorem 1.8 implies that
$\operatorname{Ldim}(M_{n})\geq\operatorname{Ldim}\left(\overline{\mathcal{H}}_{\gamma^{\prime}}^{d}\right)=\Omega\left(\sqrt{\log d(n)}\right)=\Omega\left(\sqrt{\log n}\right).\qed$
Appendix A Equivalence of Stability and List-Replicability
We prove the following more refined version of Theorem 1.2.
Theorem A.1.
Let $\mathcal{H}$ be any total or partial concept class on the domain $\mathcal{X}$. Then for every $\epsilon\in(0,1)$,
$\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{H})}\qquad\text{ and }\qquad\operatorname{LR}_{\epsilon}(\mathcal{H})\leq\frac{1}{\rho^{\mathrm{gs}}_{\epsilon/3}(\mathcal{H})}.$
Consequently, $\rho^{\mathrm{gs}}(\mathcal{H})=\frac{1}{\operatorname{LR}(\mathcal{H})}$.
Proof.
We first prove that $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{H})}$. Let $\epsilon>0$ be an accuracy parameter, and let $\bm{\mathcal{A}}$ be an $(\epsilon,L)$-list replicable learner for $\mathcal{H}$ with sample complexity $n=n(\epsilon,\delta)$. Let $\mathcal{D}$ be any realizable distribution on $\mathcal{X}\times\{\pm 1\}$, and let $h_{1},\ldots,h_{L}$ be the list of hypotheses guaranteed by Definition 1.1.
By the pigeonhole principle, at least one $h_{i}$ satisfies
$\Pr_{\mathbf{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\bm{S})=h_{i}]\geq\frac{1-\delta}{L}.$
Since this statement holds for arbitrary $\delta>0$, $\bm{\mathcal{A}}$ is itself an $(\epsilon,\rho)$-globally stable learner for all $\rho<\frac{1}{L}$. We may conclude that $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{H})}$.
Next, we prove that $\operatorname{LR}_{\epsilon}(\mathcal{H})\leq 1/\rho^{\mathrm{gs}}_{\epsilon/3}(\mathcal{H})$. Let $\epsilon>0$ be an accuracy parameter, and let $\bm{\mathcal{A}}$ be an $(\epsilon/3,\rho)$-globally stable learner for $\mathcal{H}$ with sample complexity $n_{0}=n_{0}(\epsilon)$. By the stability assumption, for every realizable distribution $\mathcal{D}$ on $\mathcal{X}\times\{\pm 1\}$, there exists $h^{*}:\mathcal{X}\to\left\{\pm 1\right\}$ satisfying
$\mathcal{L}_{\mathcal{D}}(h^{*})\leq\frac{\epsilon}{3}\ \text{ and }\ \Pr_{\begin{subarray}{c}\bm{S}\sim\mathcal{D}^{n_{0}}\end{subarray}}[\bm{\mathcal{A}}(\bm{S})=h^{*}]\geq\rho.$ (10)
For every $h\in\{\pm 1\}^{\mathcal{X}}$ and realizable distribution $\mathcal{D}$, define
$p(h)\coloneqq\Pr_{\mathbf{S}\sim\mathcal{D}^{n_{0}}}[\bm{\mathcal{A}}(\bm{S})=h],$
Denote $L\coloneqq\left\lfloor\frac{1}{\rho}\right\rfloor$, so that $\rho\in\left(\frac{1}{L+1},\frac{1}{L}\right]$, and let $\alpha\coloneqq\rho-\frac{1}{L+1}>0$. Define the list $\Lambda$ of good and likely hypotheses
$\Lambda\coloneqq\left\{h\in\left\{\pm 1\right\}^{\mathcal{X}}\ |\ p(h)>\frac{1}{L+1}\text{ and }\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon\right\}.$
Note that $|\Lambda|\leq L$ and $\Lambda$ is nonempty, as it contains $h^{*}$. Therefore, to construct an $(\epsilon,L)$-list replicable learner, it suffices to show that for any confidence parameter $\delta>0$, the learning rule outputs a hypothesis from $\Lambda$ with probability at least $1-\delta$.
Let $t\coloneqq t(\alpha,\delta)$ and $n_{1}\coloneqq n_{1}(\epsilon,t)$ be sufficiently large integers to be determined later. We propose the following learning rule $\bm{\mathcal{A}}^{\prime}$ with sample complexity $tn_{0}+n_{1}$.
Algorithm 2 The learning rule $\bm{\mathcal{A}}^{\prime}$
1:  Sample a dataset:
$\bm{S}=({\bm{P}},\bm{Q})\sim\mathcal{D}^{tn_{0}+n_{1}},\quad\text{where }{\bm{P}}=({\bm{P}}_{1},\ldots,{\bm{P}}_{t})\sim(\mathcal{D}^{n_{0}})^{t}=\mathcal{D}^{tn_{0}},\quad\text{and}\quad\bm{Q}\sim\mathcal{D}^{n_{1}}.$
2:  Define the empirical estimate of $p(h)$ as
$\operatorname{freq}_{\bm{P}}(h)\coloneqq\frac{|\left\{i\in[t]\ |\ \bm{\mathcal{A}}({\bm{P}}_{i})=h\right\}|}{t}.$
3:  Output any hypothesis $h\in\{\pm 1\}^{\mathcal{X}}$ satisfying:
  • $\operatorname{freq}_{\bm{P}}(h)\geq\rho-\tfrac{\alpha}{2}$
  • $\mathcal{L}_{{\bm{Q}}}(h)\leq\tfrac{2\epsilon}{3}$
If no such $h$ exists, output an arbitrary $h$ corresponding to “failure.”
Denote by $\mathcal{Y}$ the set of all $h$ with $\operatorname{freq}_{\bm{P}}(h)>0$ in Algorithm 2, and note that $|\mathcal{Y}|\leq t$. To show that $\bm{\mathcal{A}}^{\prime}$ outputs a hypothesis from $\Lambda$ with probability at least $1-\delta$, we will condition on the events
$A:\quad$ $|\mathcal{L}_{\mathcal{D}}(h)-\mathcal{L}_{\bm{Q}}(h)|\leq\frac{\epsilon}{3}\text{ for all }h\in\mathcal{Y}$
$B:\quad$ $|p(h)-\operatorname{freq}_{\bm{P}}(h)|<\frac{\alpha}{2}\text{ forall }h\in\left\{\pm 1\right\}^{\mathcal{X}}$
To guarantee that both events are likely, we prove the following claim.
Claim A.2.
There exist integers $t(\alpha,\delta)$ and $n_{1}(\epsilon,t)$ such that
$\Pr_{{\bm{P}}\sim\mathcal{D}^{tn_{0}}}[B]\geq 1-\frac{\delta}{2}\quad\text{and}\Pr_{{\bm{Q}}\sim\mathcal{D}^{n_{1}}}[A]\geq 1-\frac{\delta}{2}.$
Proof of A.2.
For the choice of $t$ and the proof of the first inequality, we use the uniform convergence property of the family of indicator functions on {±$\{\pm 1\}^{\mathcal{X}}$. More precisely, for $f\in\left\{\pm 1\right\}^{\mathcal{X}}$, define $\mathbb{I}_{f}:\left\{\pm 1\right\}^{\mathcal{X}}\to\left\{0,1\right\}$ as
$\mathbb{I}_{f}(f^{\prime})\coloneqq\begin{cases}1&f^{\prime}=f\\ 0&\text{otherwise}\end{cases}.$
The class
$\mathcal{I}\coloneqq\left\{\mathbb{I}_{f}\ |\ f\in\left\{\pm 1\right\}^{\mathcal{X}}\right\}$
has VC dimension $1$, and therefore, it satisfies the uniform convergence property. For ${\bm{P}}_{i}\sim\mathcal{D}^{n_{0}}$, $\bm{\mathcal{A}}({\bm{P}}_{i})$ induces a probability distribution $\mu$ on {±$\left\{\pm 1\right\}^{\mathcal{X}}$, and we have
$1-p(h)=\Pr_{{\bm{P}}_{i}\sim\mathcal{D}^{n_{0}}}[\bm{\mathcal{A}}({\bm{P}}_{i})\neq h]=\mathcal{L}_{\mu}(\mathbb{I}_{h}),$
while $1-\operatorname{freq}_{\bm{P}}(h)$ corresponds to the empirical loss of $\mathbb{I}_{h}$ on $(\mathbb{I}_{h_{1}},\ldots,\mathbb{I}_{h_{t}})\sim\mu^{t}$. Thus, by the uniform convergence property on $\mathcal{I}$, our claim holds.
Now that we have $t$, we can define $n_{1}$ and prove the second inequality. Note that for every $h\in\{\pm 1\}^{\mathcal{X}}$, for ${\bm{Q}}\sim\mathcal{D}^{n_{1}}$, $\mathcal{L}_{\bm{Q}}(h)$ is an average of ${n_{1}}$ samplings of a Bernoulli random variable with expectation $\mathcal{L}_{\mathcal{D}}(h)$. Thus, by Hoeffding’s inequality, there exists $n_{1}=n_{1}(\epsilon^{\prime},t)$ such that
$\Pr_{{\bm{Q}}\sim\mathcal{D}^{n_{1}}}\left[\left|\mathcal{L}_{\mathcal{D}}(h)-\mathcal{L}_{\bm{Q}}(h)\right|>\frac{\epsilon}{3}\right]\leq\frac{\delta}{2t}.$ (11)
Thus, by the union bound, we have
$\Pr_{{\bm{Q}}\sim\mathcal{D}^{n_{1}}}\left[\left|\mathcal{L}_{\mathcal{D}}(h)-\mathcal{L}_{\bm{Q}}(h)\right|\leq\frac{\epsilon}{3}\text{ for all }h\in\mathcal{Y}\right]\geq 1-\frac{\delta}{2}.$ (12)
A direct consequence of Claim A.2 is that
$\Pr_{\mathbf{S}\sim\mathcal{D}^{tn_{0}+n_{1}}}\left[A,B\right]\geq 1-\delta.$
Condition on events $A$ and $B$, and let $h^{*}$ be a stable hypothesis for $\bm{\mathcal{A}}$, as described in Eq. 10. We will show that $h^{*}$ is a candidate for output, so $\bm{\mathcal{A}}^{\prime}$ will not output “failure”. To check the first condition for output, we combine $B$ and Eq. 10 to show that
$\operatorname{freq}_{\bm{P}}(h^{*})\geq p(h^{*})-\frac{\alpha}{2}\geq\rho-\frac{\alpha}{2}.$
Moreover, $\rho-\frac{\alpha}{2}>0$, so $h^{*}\in\mathcal{Y}$. We may therefore apply $A$ to show that $h^{*}$ satisfies the second condition for output,
$\mathcal{L}_{\bm{Q}}(h^{*})\leq\mathcal{L}_{\mathcal{D}}(h^{*})+\frac{\epsilon}{3}\leq\frac{2\epsilon}{3}.$
Finally, let $h_{o}$ be any output of $\bm{\mathcal{A}}^{\prime}$, conditioned on $A$ and $B$. Then, $h_{o}$ satisfies the condition $\operatorname{freq}_{\bm{P}}(h_{o})\geq\rho-\frac{\alpha}{2}$, so because of $B$,
$p(h_{o})>\operatorname{freq}_{\bm{P}}(h_{o})-\frac{\alpha}{2}\geq\rho-\alpha=\frac{1}{L+1}.$
Furthermore, $h_{o}$ also satisfies the condition $\mathcal{L}_{\bm{Q}}(h_{o})\leq\frac{2\epsilon}{3}$, so because of $A$,
$\mathcal{L}_{\mathcal{D}}(h_{o})\leq\mathcal{L}_{\bm{Q}}(h_{o})+\frac{\epsilon}{3}\leq\epsilon.$
Thus, $h_{o}$ must be in $\Lambda$. ∎